OpenStudy (anonymous):

@agent0smith so if that's the a, does that mean that the answer is Logx^3(9y^2)/z?

OpenStudy (agent0smith):

Yep. \[\Large \log(x^3)(9y^2)-\log z = \log \frac{ 9x^3y^2}{ z }\] (that's the same answer, just rearranged a bit)

OpenStudy (anonymous):

ahhhhhhhhhhhhhhhhh!!!!!!! i can't believe it!! yess!!

OpenStudy (anonymous):

is the y intercept for log(3,2x+1) (0,0)??

OpenStudy (anonymous):

i didn't know how to write that, it's log, base 3, (2x+1)

OpenStudy (anonymous):

i made it equal to y.

OpenStudy (anonymous):

y=log, base 3, (2x+1)

OpenStudy (agent0smith):

\[\Large y = \log_3 (2x+1)\] To find the y intercept, plug in x=0

OpenStudy (anonymous):

a website said to bring the base over and make y an exponet so then it was 3^y=2x+1

OpenStudy (anonymous):

then I made y = 0, and any number to the 0 power is 1

OpenStudy (anonymous):

so then I had 1 =2x+1 and ended up with 0=2x, divide both sides and i have 0.

OpenStudy (anonymous):

i have never done that before so i'm hoping it's right.

OpenStudy (agent0smith):

Yes, it would be (0,0). But it's faster if you plug in x=0 \[\Large y = \log_3 (2*0+1) = \log_3 (1) = 0\]

OpenStudy (anonymous):

yyaayyy!!

OpenStudy (anonymous):

can you explain that a little? how did you get 0 that way?

OpenStudy (agent0smith):

Log of 1 is equal to zero, no matter the base.

OpenStudy (anonymous):

ohhhhh...cool.

OpenStudy (anonymous):

what is Find f^-1(x) for f(x)=6x-5? do i just make it y=(6x-5)/-1

OpenStudy (agent0smith):

That means you have to find the inverse function. y = 6x-5 swap the x's and y's... x = 6y - 5 now rearrange it till you get y=... That'll be the inverse, the f^-1(x)

OpenStudy (anonymous):

ohhhhhh....

OpenStudy (anonymous):

so simple